OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x − 3y = 3 5x − 4y = 3 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system −3 3

mathslover (mathslover):

4x-3y = 3 5x -4y =3 this means that 4x-3y = 5x-4y 5x-4x=-3y+4y x = y

mathslover (mathslover):

so we have 4x-3x = 3 x = 3 and similarly y = 3

mathslover (mathslover):

got it @ConnorYoung89

OpenStudy (anonymous):

Thank you so much.